Wednesday, 28 January 2026

Wednesday 28th January 2026

Mike Kushy, Stu Jellis, Tom Frost and Jake Baynes. First trip back in since closure and the last trip in December 3rd 2025, so personally I was very keen to see the conditions in MHH. Fair stream heading into Puddle Dig as we made our way in. Everyone got through the light connection fine, it was TF's first trip, JB's and SJ's third time. Comments were made about the significant improvments to the Light Connection and The Hookah Pipe to get all the team to the end (previously JB was unable to get through The Hookah, this time he went head first!). Unfortunately MHH was as much a lake as we've ever seen it, although the water has been tracing it's way to the far corner and the amount of water coming in today didn't warrant the lake size (it should having been overflowing back into TPC at that rate!). We did a bit of poking in the mud then headed out. On the way out we popped into 'ED's bet'. SJ went for a look first and made it to where Ed got before. MK also went in for a look and got to the same place. He had brought the long crowbar back from TPC and poked at the floor a bit, right next to the upright flake boulder. Stones did freefall below for an estimated 15ft, it has promise but stabilisation is a must and personally I'd use remote rock removal techniques to push further (video was made). MK and TF went out the old route, whilst the others headed out via Puddle dig. SJ investigated Puddle dig and reported no water at all, just mud, although there was a healthy stream running in! The cave seemed to be breathing the whole trip, it was a colder evening outside with intermittent showers. 2 hours. Mike.

Video of Ed's bet.

SJ, JB and TF in MHH lake. Photo MK.

Friday, 16 January 2026

Friday 16th January 2026

White Rabbit is back open. 

We've been asked to be discreet when changing (put on fleece before arriving or just be discreet).

Also keep noise to minimum and lights low, the people who are in the farmhouse opposite where we park the cars are not the same as those who own White Rabbit.

Please pull the gate to when digging and after leaving. We mean the second gate that is on the cattle grid (basically keep the WRH parking secure when were underground and leaving).

Sunday, 7 December 2025

Sunday 7th December 2025

**

Ownership of the land has changed and we are asking about parking and other access arrangements.

We should not drive up, park or go across to the dig at all until we know more.

Hopefully update soon.


**

Thursday, 4 December 2025

Wednesday 3rd December 2025

 Eds Bet

Ed Ford and Mike Wise entered WR at 1830. We took 2 drills, batteries, plug and feathers and a crowbar. Got to Eds Bet which looked to be fairly sketching digging. EF started by removing some sizable blocks from the floor working above 2 larger blocks. Some of the removed blocks had holes on the sides from the 12g bang of the Up&Over Boulder - one of the bigger ones brutally attacked one of MWs fingers and crushed it, chipping the top joint, whilst being wrestled into its storage location although the extent of the damage wasn't appreciated until the next morning 😢.

Once a bit of loose fill had been removed, it was clear there was a drop below taking some time for loose rocks to fall and a flowstone covered vertical fault wall was visible.

 EF asked MW to take a look to ensure he didn't bring the roof down on him. MW then took a turn removing some of the spoil and a few loose rocks threatening to drop on anyone venturing below the the 2 larger blocks. The drop below became more visible but all looked reasonably stable considering.

We swapped again and EF went below the two larger blocks. Down here there is sizable space between the bedding. Some overly large rocks were lifted out before the whole floor became quite unstable. Careful footing was required at this point and MW requested that EF refrain from getting sucked down into a loose rock funnel as it was nearing pub time. All the rock was removed bar one which was far bigger than could be lifted. To this point we hadn't used the P&F and decided we didn't have time to retrieve the hammer from Tea Party. EF instead opted to remove the keystone holding it and drop it. It dropped a couple of inches onto a shelf also exposing a body sized hole. EF headed through the hole to explore the new bit before heading out reaching the surface at 2100 for a celebratory pint and a cheese&onion roll. The new bit looks promising and is not fully explored - we will be back. It is still a 2 caver job.

Ed Ford and Mike Wise

 Edit - in the initial rock removal EF found an artefact, that despite its resemblance to the plug of a 14mm plug & feather set, we are convinced is a bone from the mysterious and long extinct Sabre Toothed White Rabbit.  We left it in situ pending professional examination. 

 Photo by S.Wise taken next morning. 

Mikes finger after being crushed by a large rock. Its twice the size it should be.


 

Friday, 21 November 2025

Friday 21st November 2025

Mike Kushy, Hollie London, Alessio Sancetta, Sam Chung, Tuska, Wayne Starsmore, John Cooper and Brian Prewer. We all met around 2.30 and quickly went over how to set up the radio location equipment. Then MK, HL, AS and SC went to the first chamber on the old route in, just to check comms. Comms was established but too close to the iron ladders to Radio Locate. We then all headed to TPC to set up. SC and AS worked on this whilst MK and HL cleared The Hookah Pipe of the bang spoil and tried to divert puddles. Once TPC was located HL, AS and SC went on to MHH...I also believe the team carried on digging out channels to the far corner and the water reaches there now. Meanwhile MK made 2 mixes and carried on with supporting the landing pad. The MHH location sounded really successful (although comms weren't great), now to cross reference with the survey. 1 drum of sand, 1 of cement, lots of stone dust and water in cave. Mike.

AS and JC setting up on surface. Photo SC.

WS, AS, HL, T and MK. Photo SC.

AS on comms in TPC. Photo SC.

HL and AS in MHH. Photo SC.

The Hookah Pipe after clearing. Photo MK.

Friday, 14 November 2025

Friday 14th November 2025

Mike Kushy and Sam Hill. Possibly the wettest day of the year so far with 1 1/2 inch of rain in Priddy within 24 hours, fairly warm though. We took in 2 drills, HE and a full size spade. The entrance shaft was wet and loads of inlets higher up than we've ever seen before. Surprisingly the Lower Bedding and Cement Dig had no drips at all, whereas this used to be the wettest area. SH used his drill to put an anchor in for the Light Connection rope (need to take mallion in for this), whilst MK set to laying in a puddle drilling shot holes in The Hookah Pipe. Really apparent and cold draught coming out from MHH today. SH then went into MHH with the spade for some mud mining, he also took some footage. MK then finished holes and joined SH where we started digging away mud from the far corner. Massive lake in the chamber, but not as much water coming in as I thought there would be. MK laid the charge with SH assisting. On the way out SH went to check out Puddle Dig and reported there's a clear puddle but it's not as deep as it should be considering the amount of water coming in! All fired fine. 2 1/2 hours. Mike.

Video of MHH in flood, by SH.

MK in MHH. Photo SH

Wednesday, 12 November 2025

Wednesday 12th November 2025

Ed Ford and Mark Edwards went down 1815. One cement and one sand taken down. Hookah Pipe cleared of bang debris. A torrent of water was coming down in Mad Hatters Hall with the current dig flooded. This is definitely taking the water though as it flows into it. No problem with air. Felt fresher in Mad Hatters Hall than rest of cave! Out by 2030. Ed.

Saturday, 8 November 2025

Saturday 8th November 2025

Party including Mike Kushy and Tuska fitted the extra bolts into the gate hinges. Mike.

Thursday, 30 October 2025

Wednesday October 29th

Mike Wise and Ed Ford. Our mission was to clear the debris from last week's Hookah Pipe (HP) diameter adjustment and if needed repeat the process. We declined the suggestion to take in more cement so just took a 36v drill, a couple of 10mm drill bits, a spare set of batteries and passage widening kit in two bags.

A slight whiff of fumes was detected from the pit above the Light Connection (LC). MW descended to check things out while EF waited above. It seemed OK at first; there was a pile of small debris in the HP but further slight adjustment was needed and so the drill bag came down. There was no detectable draught and TPC was the same temperature as the upper cave. But then the fumes became more apparent in the area around HP and upstairs and the drill bag returned up LC. EF came down for a look and chat.  

EF declined to enter HP; his sense of smell hasn't been wrecked by time like mine has. We did discuss moving some rocks towards the platform below LC keeping us away from the lower part of TPC, but then EF traversing the rock pile found a pocket of fume. Time was called and we returned up LC. Despite using exemplary thrutching technique we were both panting on reaching the top and according to EF, MW was sweating buckets.

We exited puffing and panting with throats complaining we had just eaten a handful of gravel. No photos taken. Medicinal substances were taken in The Hunters' which helped to improve things slightly.

1 hour 40 minutes

 Mike Wise

Monday, 27 October 2025

Monday 27th October 2025

Mike Kushy and Tuska. 2 hours fitting metal gate. Need an extension bar for the driver to get the lower screws in on the hinges and need to sort out a lock. Mike.

Gate fitted. Photo MK.

Gate fitted. Photo MK.

Sunday, 26 October 2025

Sunday 26 October 2025

Mike Kushy and Tuska. MK picked up a piece of 3mm sheet metal that John Gisborne had cut for us, thanks JG! The sheet was then taken to Tuska's house where we cut to hole size and attached hinges and handles. Gate now ready to fit! 2 hours. Mike.

Top side of Gate. Photo MK.

Underside of Gate. Photo MK.

Wednesday, 22 October 2025

Wednesday 22nd October 2025

Mike Kushy, Mike Wise and Ed Ford. We met up at 4pm and headed in together. MW and EF went in the usual route carrying bags whilst MK wound up and collected the top wire.,Tthis was taken further down to leap frog the lower wire. MW set to drilling and removed the 1m drill bit so MK could go ahead and grab the newer rope (the air felt quite thick in the intermediate chamber and drill dust was just hovering). Leaving MW to drill shot holes MK and EF replaced the rope on the light connection and set about building the landing area properly with rocks and mortar. 3 lots of sand left in cave and half a cement, loads of water. MW then used what battery was remaining to drill a bolt hole to create a traverse type line (this is only half drilled). MK meanwhile set the charge. We had a little look down 'Ed's bet' (the vertical pot directly below Cement Dig) and it does look worth going in and clearing the rocks out at some point soon. We fired from the top of the Lower Bedding, heard and felt faintly. 3 hours. Mike K.

MW and EF building in TPC. Photo MK.

MW and EF building in TPC. Photo MK.

Sunday, 19 October 2025

Saturday 18th October 2025

Mike Kushy, Hollie London, Mikey Waterworth and Olivia Dawson. MK and HL to begin with started at 10.30am, we cleared the bang spoil from the up and over boulder to start. The space over the top is a lot more comfortable, although the lower piece is also shattered and could be removed entirely (personally I think we should just cement in and stabilise what's there). Next job was to drill the stemples in (thanks to Lee V for use of your Hilti drill for this), one on the side wall below the ledge and another up near the scaff' bar. We could do with putting a couple of smaller metal holds higher up again, but all a lot safer now. Next MK started drilling for shot holes in The Hookah Tube, 3 holes @600 and 1 at just shy of 1m (1m drill bit currently sticking out of the wall, battery ran out!), whilst HL went on into MHH and started to dig out the final squeeze some more. A retaining wall was built at the 'up cave' end of MHH to put all the glorious mud behind. CO2 and O2 tests were done throughout, see results below. Around 12.30 MW and OD came and met us in MHH. They said the air looked very thick. We agreed a plan and continued the already started dig to the right of the formation. Unfortunately this hit a solid wall 1m in and won't give us the clearance to dig around the formation so we will have to try a dig in the far corner? (tarpaulin sheet now over the formation for protection). MK and HL left the others to continue digging as they both felt puffy and like they needed better air. Really felt the air quality change once back in TPC. HL started to rebuild the landing area below the Light Connection. We had a tidy round and got packed to leave around 3pm. MK went through to MHH to let the others know our plan and we all decided to all leave together (also getting close to 3% CO2, our set safety limit). Very wet in MHH today with puddles everywhere. Would recommend not going into MHH for a while and let the air settle, so we can retest the differences (probably the worst day of the year today for testing air quality in a cave, so hopefully it passes). Oh MK and HL went to mix some muck towards the end but realised there was no cement in the cave. About 5 drums of sand though, 10l of mixing water and a bag of stone dust. 5 hours. Mike K. 

MK drilling holes for stemples. Photo HL.

CO2 and O2 results (thanks to Stuart 'Mac' for use of your monitors):

Start (11.30am) - CO2: 2.08, O2: 18.76%

12.01pm - CO2: 2.31, O2: 18.5%

End (14:16) - CO2: 2.56, O2: 18.18%

*Apart from 15 minutes (around 1pm) there was only 2 people in MHH at any time and working whilst we monitored*

Friday, 17 October 2025

Friday 17th October 2025

Hollie London, Harley Bennett

HL and HB headed down WR at 1630 to assess the damage from the previous bang. The new ladder made an excellent addition to the entrance series and new tape definitely highlights where you need to go navigationally. HB highlighted the bacon as we entered the bedding chamber, something we would normally charge past on the way to the dig face. After stopping to photograph the bacon, we continued down to the up and over boulder. The bang has successfully doubled the body space above the boulder, with the bottom half remaining intact. The pot before the boulder looks to have opened up and was drafting well. Due to lack of man power and everything being lose, we decided not to clear and retreated having photographed the damage. On exit, one plank fell off the wooden lid down the  entrance and needs a repair job to screw down. HL made a short term fix until the weekend so nothing can fall down. Hollie.

New base section. Photo HL.

Formations in Lower Bedding Chamber. Photo HL.

Needs clearing. Photo HL.

Wednesday, 24 September 2025

Wednesday 24th September 2025

Mike Kushy, Mike Wise and Sam Hill. All down the cave for around 6.30pm. Took in drill and extra batteries plus charge materials. The wire was tested from the shaft as it hadn't been used for a while, no signal so took the tester to the joint. Tested higher wire from lower end and all fine, tested lower wire and no signal. Took the tester to the furthest end of the wire and discovered stripped cable and ends broken, fixed and tested okay. MW began drilling the up and over boulder, whilst SH cleared out behind where the recent rock moved ready for cementing. MK made some mixes in TPC and sent them up to SH. In the meantime MK kept moving rock around in TPC to make the floor better. Around 8.30pm MW called that he was all set up so we decided it was a good time to leave. Charge set off from Boulder Rift Chamber, very quiet considering we were only 30m of bang wire away. 2 1/2 hours. Mike K.

Tuesday, 23 September 2025

Tuesday 23rd September 2025

John Cooper and Mike Kousiounis (Kushy). Survey trip of Tea Party Chamber and on to current end. Twenty minutes in, 60 minutes surveying and 40 minutes out (it’s uphill!). The current end chamber appears to be developing a wetter, muddier floor. Now got to process the data collected. (See Survey Page.) 2 hours. John.

Saturday, 20 September 2025

Saturday 20th September 2025

Mike Kushy, John Arnold (With GoPro), Jason Hares (Jay) (With GoPro) and Lucas McMaster.  All undergound just after 10am. Took in sand, stemples and fixing kit, 5l of mixing water and food. MK went ahead using pink conservation tape to set route markings (don't cross the tape for the best route basically).  All down to TPC no problems, JH mentioned that the rock above the light connection would need sorting on the way out. MK sorted out more taping in TPC whilst the others took the time to look around. We then all went through to the final chamber where more taping was done by MK whilst the others used GoPros on sticks on various holes. MK then went through with the others on how to clear the 'intermediate chamber' and left them to it. The plan was to go and set the via ferrata stemples in place. Unfortunately the drill started smoking, so that plan didn't work. With JH in TPC hauling the buckets, MK part way in assisting buckets through the pipe, LM filling buckets/working the rope and JA loosening mud we cleared for a good hour before stopping for lunch. After lunch MK went to loosen mud and had discovered JA had been digging his way to China. With MK loosening mud and digging to China, JA filling buckets/working rope and LM/JH hauling back to TPC we did another good hour or so more clearing. Towards the end we discovered a hard calcite floor about 1ft deep, hopefully all the mud is only 1ft deep! Packed away drill and we all started to head out, also with any empty bags. The boulder above the light connection did need sorting indeed, LM and JH sorted this out and got it moved close to the 'emergency exit'. Consequently some clearing of loose stuff and stabilising needs doing, this is a high priority job. Need to make a list of materials in TPC now. 6 hours. Mike K.

The link to the video is here: https://www.youtube.com/watch?v=JyboaDlXhcg

It's rather long so key bits can be found here:

0.19 Cement Dig
0.23 Approaching Light Connection
0.35 All in Tea Party Chamber
0.45 Entering The Hookah Pipe
0.55 to 1.09 In the Final Chamber - note the frost shattered Stal
1.24 Break in filming - starts again leaving Light Connection
1.38 Rock moves!

Wednesday, 17 September 2025

Wednesday 17th September

Sam Hill, Mikey Waterworth, Ed Ford, Jake Baynes, Adrian Miles and Mike Wise.

Team split into 2 parties with different aims

Mike Wise solo report:

A drill, capping kit, lump hammer and micro crowbar were taken down to Boulder Chamber. Here the irritating pyramid shaped lump on top of the large boulder at the top of the 10ft climb was capped off using 3 holes and 5 caps. The top of the boulder is now flat. Some minor dressing was done with the lump hammer. The kit was abandoned temporarily in the chamber.

The hammer and crowbar were taken down to Boulder Rift Chamber where the floor of the step-up and crawl into the bedding below Calcite Squeeze was attacked for about 45 minutes to improve head room and generally remove lumps from the floor. Large amounts of rock varying in size from very big/just liftable to hand size were removed along with their covering of mud. The big stuff was stacked in the chamber, the smaller stuff and some mud was rammed back into the gaps that appeared below the mud covering.

Started packing up at 2045 and began slowly exiting at 2100 having heard nothing from the team below. They caught up with me at Boulder Chamber and at the entrance shaft I found that someone was jiggling the hauling rope from the top. It turned out to be SH who mistakenly had exited via the old route - he must have run out!  Thanks to JB for carrying my drill box down into Boulder Chamber at the start. 

Other report:

All visited TPC. EF's foot came off the slippery scaffold bar on the descent from the Light Connection and he somersaulted down onto the chamber floor. Injuries declared in the Hunters later were a damaged knee and a badly swollen wrist. MWa scanned everything from TPC to the end.

2 hours 

Mike Wise

After Ed’s acrobatics, myself and Jake had a look at TPC. I then went through the Hookah to join SH and MWa at the end. Myself and MWa went back to the chamber just after the hookah and cleared some of the spoil in both directions. MWa and SH went back to the final chamber to look at prospective leads whilst the rest of us started heading out. All on the surface by 21:30. Adrian. 

Sunday, 14 September 2025

Sunday 14th September 2025

Mike Kushy, Ed Ford and Hollie London. MK and EF to begin with. Headed in just after 10am leaving one of the new 6l daren drums for HL to pack her camera kit in (thanks to Rob Harper for the kind donation). EF took in his drill and picked up the drinking water on the way in and MK took charged batteries, drum of sand and a 5l container of mixing water in. When we got to TPC EF got to making some brews, then HL arrived equipped with her camera and some hauling rope (which proved very useful later). After morning tea we had a go at The Hookah Pipe some more, with MK on the far side pf'ing back and EF/HL working the TPC side. When we decided it looked big enough to haul buckets through we all stopped for lunch. EF and HL then went through to the most recent chamber and we used the haul line to pull buckets back. Ideally we need flat laid 5l buckets cut to 3/4's for this, plus we can still open the whole tube up some more (maybe a small charge is needed on one side). MK pulling the line in TPC and EF/HL taking turns pushing on, 66 buckets in total. Around 59 buckets MK shouted to call it at 60 but the reply was 'There's a great big chamber not far off'. All 3 now in the smallish chamber we all took turns loosening mud and fishing out rocks. First one through then stopped and cleared from the otherside for others, all got through in the end. The new new chamber is roughly 10m long x 4m wide x 3m height with nice broken column formation at the far end. Lots of rifts intersecting but the draught is not obvious anymore. Liquid mud floor which is very shiny (perhaps a small lead content?). We will have to discuss the next steps as a team once everyone has had a look. 4 drums of sand in cave, 1 drum of cement, 5l of mixing water and 10l of drinking water. 6 1/2 hours. Mike

Click here for todays video.

HL coming back out of the newest chamber. Photo MK.

Fossil in new chamber. Photo HL.

EF in the new chamber. Photo HL.

MK looking at the broken column formation. Photo HL.

EF entering the new chamber. Photo HL.

Same photo as previous but further back. Photo HL.

EF fishing out rocks from the soon to be breakthrough point. Photo HL.

The column found in the 'intermediate' chamber. Photo HL.

Wednesday, 10 September 2025

Wednesday 10th September 2025

Mike Kushy, Hollie London, Ed Ford and Mike Wise. MK and HL to begin with. We made our way in around 5.30pm with batteries and a tub of sand. Picking up 2x scaff' bars on the way everything was taken to TPC. EF arrived not long after with his Dewalt 18v drill so we handed him some pfs and he opened up the head of the light connection more. This has revealed that some extra stabilisation needs doing here, possibly some cementing and a scaff' bar for a brace. Meanwhile HL was attacking The Hookah Pipe with the sds and chisels, whilst MK did more furnishing of TPC's floor. MK then went through The Hookah Pipe and used the Bosch sds from the far side. MW arrived and made it into TPC for the first time and said it wasn't as bad as we made out. He brought his Makita 36v in and attacked the pipe from TPC side. We all swapped turns tidying the floor and chiselling. Towards the end EF managed to get through THP into the decorated chamber beyond so he did some pf work from the far side. HL had a go at THP too. I think she could have done it but the squeeze is very intimidating and it's not going anywhere. Probably another session and we may be able to start hauling buckets of mud back into TPC. 3 tubs of sand and 1 tub of cement in cave, no water. There is 5l bottle of drinking water on the shelf near the bottom of the shaft. This will need an empty tackle sack to take further in. We also need to take in a heavier lump hammer for calcite.  3 1/2 hours. Mike K.